You are given an unknown metal. You obtain the following data for the unknown metal: The mass is 3.50kg and when 1.319 x 10^4J of heat is added to this sample, (initially at a temperature of 21.33 C) the temperature increases to 48.25C. Identify the unknown metal.

Metal                                             Specific Heat (J/g C)

Aluminum                                      0.902

Copper                                          0.385

Gold                                               0.129

Iron                                                0.450

Zinc                                               0.140

Lead                                              0.132
Answer #1

Change in temperature ∆T = 48.25-21.33 = 26.92C,

Mass = 3.5kg = 3500g

Heat = specific heat*mass of metal*∆T

1.319*10^4 = specific heat*3500*26.92

Specific heat = 1.319*10^4/94220 = 0.1399 = 0.14 J/g

So, the metal is lead (Pb).